773 mK/s = _______ K/s

773 mK/s = _______ K/s
Answer #1

773 mK/s

1mill= 10-3 = 0.001

773 mK/s = 773 x 0.001K/s = 0.773 K/s

so, 773mK/s = 0.773 K/s